OpenStudy (anonymous):

Use the Quadratic Formula to solve the equation. -2x2 - 5x + 5 = 0

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[ \\ \sqrt{b^2-4ac}=...?\]

OpenStudy (anonymous):

@hartnn but that is for finding the discriminant. But to solve the equation, you have to use the formula \[-b \pm \sqrt{b^2-4ac} /2a\]

hartnn (hartnn):

yes, that is the next step if you notice, the formula has sqrt(b^2 -4ac) so we better find it before using the formula :)

OpenStudy (anonymous):

ok now what?

hartnn (hartnn):

what values did u get ? for a,b,c ?

OpenStudy (anonymous):

idk how am i suppose to plug them in? My computer is being dumb saying "math processing error"

hartnn (hartnn):

-2x^2 - 5x + 5 = 0 compare it with ax^2+bx+c =0

OpenStudy (anonymous):

so a is -2? b is -5 and c is 5? :)

hartnn (hartnn):

thats correct! now calculate sqrt (b^2 -4ac) = ...

OpenStudy (anonymous):

(-5^2-4(-2*5)) ?

hartnn (hartnn):

careful with brackets \((-5)^2 - 4 (-2)(5) = ... ?\)

OpenStudy (anonymous):

So its not like that ^^?

hartnn (hartnn):

oh sorry, i forgot... (-5)^2 - 4 (-2)(5) = ...

OpenStudy (anonymous):

-5^2 = -25 right? so -25 -4 = -29?

hartnn (hartnn):

b = -5 and b^2 will be (-5)^2 = -5*-5 = +25

OpenStudy (anonymous):

oh ...im really bad at this

OpenStudy (anonymous):

My answer is suppose to be like a fraction let me show you answer choices \

hartnn (hartnn):

4 (-2)(5) = - 40 +25 - (-40) = +25 + 40 = +65 b^2 -4ac = +65 sqrt(b^2-4ac) = sqrt (65)

OpenStudy (anonymous):

So it A or D?

OpenStudy (anonymous):

Or is my answer D?

hartnn (hartnn):